# When LLM Judges Inflate Scores: Exploring Overrating in Relevance Assessment

### v2

@@ 3.2. Controlled Lexical–Semantic Variations
Prior work shows that query terms can cause LLMs to overrate non-relevant passages (Alaofi et al., 2024). To disentangle lexical overlap from semantic relevance, we construct three controlled, query-dependent variants —  Semantic-only (SEM), Lexical-only (LEX), and Query-only (QRY) — by inserting a single sentence at the beginning of passages originally labeled non-relevant by both human assessors and LLMs, where injection has the strongest influence on relevance judgments (Tamber and Lin, 2025).
@@
For the LEX variant, we construct a semantically non-relevant sentence that retains all query terms (excluding stopwords). To strictly control unigram overlap and eliminate unintended changes to the passage meaning, we use a fixed, content-neutral template: “The terms $w_{1}$, $w_{2}$, …, $w_{n}$ are shown together as an example of word grouping in text data.” where $w_{k}$ represent the $k$ individual query terms.
In the QRY variant, the inserted sentence is the original query string itself (Alaofi et al., 2024). This design aims to maximize phrase-level lexical overlap with the non-relevant passage without introducing additional semantic content beyond the query.
@@ Results.
Figure 2 reports the label distributions after inserting each variants. Under the SEM condition, the majority of the models assign 1–2 relevance labels, indicating systematic under-recognition of semantic relevance in the absence of lexical cues. In LEX condition, unigram-level lexical overlap alone is sufficient to substantially elevate relevance labels despite the absence of semantic alignment. This reveals a strong sensitivity to term-level lexical cues. The QRY condition yields severe overrating: inserting only query string frequently triggers  perfectly relevant labels for non-relevant passages which confirms prior findings (Alaofi et al., 2024). This demonstrating that phrase-level lexical anchoring can override semantic judgment. Overall, even with semantic understanding, our experiments show that LLM relevance assessment is disproportionately influenced by lexical cues, confirming strong lexical anchoring effects across models.
@@ 4. Conclusion
In this study, we investigated label overrating and model confidence in LLM-based relevance judgments across multiple open-weight LLMs and evaluation paradigms. Across all evaluation settings, we showed that overrating is a system-wide bias, particularly in graded and pairwise evaluations. Overrating is less pronounced in binary relevance judgments, suggesting that LLMs may be better suited for coarse-grained tasks that focus on distinguishing relevant from non-relevant documents. Through experiments on semantically equivalent passage variations, LLMs are demonstrated to be highly sensitive to passage length even when semantic content is preserved, while showing comparatively limited sensitivity to syntactic form. Controlled query-related variant experiments indicate that, despite some semantic understanding, LLMs remain strongly influenced by lexical cues, especially direct query-string overlap. These findings underscore the need for diagnostic evaluation frameworks that explicitly probe such biases when applying LLMs for relevance assessment.

### v3

@@ 3.2. Controlled Lexical–Semantic Variations
QID: 156493
QUERY: do goldfish grow
SEM: These orange cyprinids continue to increase in size throughout their lifespan under suitable conditions. LEX: The terms goldfish and grow are shown together as an example of word grouping in text data. QRY: do goldfish grow.
Figure 2. SEM, LEX, and QRY variants for the query “do goldfish grow”. SEM preserves relevance without surface query terms; LEX provides the query terms such that they are not relevant to the information need; and QRY simply adds the query itself.
Prior work shows that query terms can cause LLMs to overrate non-relevant passages (Alaofi et al., 2024). To disentangle lexical overlap from semantic relevance, we construct three controlled, query-dependent variants, including Semantic-only (SEM), Lexical-only (LEX), and Query-only (QRY), by inserting a single sentence at the beginning of passages originally labeled non-relevant by both human assessors and LLMs, where injection has the strongest influence on relevance judgments (Tamber and Lin, 2025). Figure 2 illustrates these three variants for the example query “do goldfish grow”, where each injected sentence is prepended to non-relevant passages to re-assess their relevance.
@@
For the LEX variant, we construct a semantically non-relevant sentence that retains all query terms (excluding stopwords). To strictly control unigram overlap and eliminate unintended changes to the passage meaning, we use a fixed, content-neutral template: “The terms $w_{1}$, $w_{2}$, …, $w_{n}$ are shown together as an example of word grouping in text data.”, where $w_{k}$ represent the $k$ individual query terms.
In the QRY variant, the inserted sentence is the original query string itself (Alaofi et al., 2024). This design aims to maximize full-query lexical overlap with the non-relevant passage without introducing additional semantic content beyond the query.
@@ Results.
Figure 3 reports the label distributions after inserting each variant. Under the SEM condition, the majority of the models assign 1–2 relevance labels, with perfectly relevant (label 3) rarely assigned, indicating systematic under-recognition of semantic relevance in the absence of lexical cues. Under LEX condition, unigram-level lexical overlap alone is sufficient to elevate relevance labels despite the absence of semantic alignment. This reveals a strong sensitivity to term-level lexical cues. While some models (e.g., Gemma) assign labels comparable to SEM, others (e.g., Qwen) remain largely at label 0, suggesting that query terms are insufficient to override semantic non-relevance for stronger models. The QRY condition yields severe overrating: inserting only query string frequently triggers  perfectly relevant labels for non-relevant passages which confirms prior findings (Alaofi et al., 2024). Gemma assigns perfectly relevant to the majority of non-relevant passages once the full query is inserted, while others (Mistral, Qwen) remain relatively robust. This demonstrates that full-query lexical anchoring can override semantic judgment in most cases. Overall, even with semantic understanding, our experiments show that LLM relevance assessment is disproportionately influenced by lexical cues, confirming strong lexical anchoring effects across models.
@@ 4. Conclusion
In this study, we investigated label overrating and model confidence in LLM-based relevance judgments across multiple open-weight LLMs and evaluation paradigms. Across all evaluation settings, we showed that overrating is a system-wide bias, particularly in graded and pairwise evaluations. Overrating is less pronounced in binary relevance judgments, suggesting that LLMs may be better suited for coarse-grained tasks that focus on distinguishing relevant from non-relevant passages. Through experiments on semantically equivalent passage variations, we demonstrate that LLM judges are sensitive to passage length even when semantic content is preserved, while showing comparatively limited sensitivity to syntactic form. Controlled query-related variant experiments indicate that, despite some semantic understanding, LLMs remain strongly influenced by lexical cues, especially direct query-string overlap. These findings underscore the need for diagnostic evaluation frameworks that explicitly probe such biases when applying LLMs for relevance assessment.
5. Future Works
The experimental setup relies only on TREC DL datasets, which primarily consist of short queries paired with relatively short passages. Future work will extend our analysis to investigate how LLM-based relevance judges behave in collections, such as TREC-8 (Hawking et al., 2000) and TREC Robust04 (Voorhees, 2006), contain substantially longer, multi-topic documents. Another concern is our current study adopts the UMBRELA prompt as a representative instantiation of LLM-based relevance judges. Prior work demonstrates that variations in prompt structure can substantially influence LLM judgment behaviors. Extending beyond a single prompt is therefore essential to establish the robustness of our findings. Future work will construct analysis on multiple prompts spanning reasoning strategies, prompt structures, and persona. These extensions will strengthen the generalizability of our findings and position the proposed analysis as a framework for studying LLM judges, rather than a result tied to a single prompt instantiation.
We acknowledge the concern that modern LLMs may have been exposed to MS MARCO (Bajaj et al., 2018) or TREC Deep Learning (Craswell et al., 2020, 2021) qrels during pre-training (Thomas et al., 2024; Faggioli et al., 2023), which could inflate apparent agreement with ground-truth labels through memorization rather than genuine judgment capability. Our experimental design partially mitigate this risk. Specifically, the summarization and expansion rewriting generates passages that are unlikely to have appeared in pre-training, and the SEM, LEX, and QRY variants are synthetically constructed rather than directly drawn from the original collection. Hence, the bias patterns we report in this work cannot be fully explained by memorization of specific query-passage pairs. However, we cannot fully rule out that LLMs retain topical or stylistic priors from MS MARCO that bias their judgments in ways not captured by our controlled perturbations. A rigorous contamination audit remains an important future work.
Acknowledgements.
We thank the anonymous referees for their insightful comments and suggestions. This work was supported by the UQ HERA initiative and by a Google Research Scholar grant.